# Dissipationless dynamics of spin supersolid states in a spin-1/2 triangular antiferromagnet with impurities

### v1

@@ III Addition results of the dynamical structure factor
As suggested by the study of the easy-axis anisotropic Heisenberg model on the triangular lattice [87], the rotonlike minimum might also appear at the $Y_{1}$ points. Here we keep the same impurity density level of $1.85\%$ and other parameters as the Figs. 3 and 4 in the main text. For the “Y” supersolid state with impurities, we use a shorter $\tau_{tot}=42/J$ for better numerical convergence. Figure S3 shows the $\chi(\textbf{q},\omega)$ at the $Y_{1}$ points between the $A$ and $B$ points for various states. As shown in Figs. S3(a) and (d), we identify a rotonlike minimum in the “Y” supersolid state which remains almost the same in the presence of impurities. This minimum at the $Y_{1}$ points is clearly visible as compared to the state without magnetic field [49], and cannot be captured by the linear spin wave theory as magnon dispersions are strongly renormalized. In the “V” supersolid state the rotonlike minimum at the $Y_{1}$ points disappears, as shown in Figs. S3(c) and (f), and the lowest energy excitations can be qualitatively reproduced by the linear spin wave results. For the spin supersolid states, the whole spectrum remains similar with impurities except for some broadening of the spectral weight. However, the splitting of the lowest band in the UUD state due to impurities can be seen by comparing Figs. S3(b) and (e).

### v2

@@ III Addition results of the dynamical structure factor
As suggested by the study of the easy-axis anisotropic Heisenberg model on the triangular lattice [89], the rotonlike minimum might also appear at the $Y_{1}$ points. Here we keep the same impurity density level of $1.85\%$ and other parameters as the Figs. 3 and 4 in the main text. For the “Y” supersolid state with impurities, we use a shorter $\tau_{tot}=42/J$ for better numerical convergence. Figure S3 shows the $\chi(\textbf{q},\omega)$ at the $Y_{1}$ points between the $A$ and $B$ points for various states. As shown in Figs. S3(a) and (d), we identify a rotonlike minimum in the “Y” supersolid state which remains almost the same in the presence of impurities. This minimum at the $Y_{1}$ points is clearly visible as compared to the state without magnetic field [49], and cannot be captured by the linear spin wave theory as magnon dispersions are strongly renormalized. In the “V” supersolid state the rotonlike minimum at the $Y_{1}$ points disappears, as shown in Figs. S3(c) and (f), and the lowest energy excitations can be qualitatively reproduced by the linear spin wave results. For the spin supersolid states, the whole spectrum remains similar with impurities except for some broadening of the spectral weight. However, the splitting of the lowest band at $B$ points in the UUD state due to impurities can be seen by comparing Figs. S3(b) and (e).